# Training on the Test Task Confounds Evaluation and Emergence

### v1

@@ 7 Discussion
The anecdote holds a lesson for the evaluation of large language models half a century later. Knowledge about the evaluation conditions necessarily influences training practices under competitive pressure. It may be a fool’s errand to prohibit the practice. Instead, we propose to adjust for it by giving every model the same task-specific preparation before evaluation. We work from the assumption that training on the test task, in general, cannot be effectively detected, disallowed, or disincentivized. Detecting what training data a model has seen is a notoriously difficult problem—existing heuristics achieve partial success at best. Researchers routinely acknowledge the futility of fighting data contamination. Moreover, we anticipate that the ways to effectively train on the test task will only grow in scope and adoption.
Our work demonstrates that comparisons of different models are confounded by the choice of training data and training practices. Different model families vary in the degree that they were—implicitly or explicitly—trained on various test tasks. It therefore makes little sense to compare model performance at face value without accounting for how the training data relate to the test task. The same problem extends to scaling. Smaller models can appear unexpectedly performant if they were trained to a greater extent on task data.
We can apply the same principles to emergent behavior. After training on the test task, model capabilities become predictable at smaller model size and grow continuously with scale. This is not to say that emergence isn’t real; it may well be a real phenomenon for a fixed choice of dataset and evaluation metric. But training on the test task removes the unpredictability and discontinuity associated with emergence, notably without any change in the metric, thus largely disarming the ominous nature of emergence.
Despite the daunting challenges that training on the test task poses for the fair evaluation of language models, it’s also its own best remedy. Giving each model the same sufficient task-specific fine-tuning harmonizes model comparisons, deconfounds scaling laws, and linearizes the relationship between model capabilities and log-scale. We hope that our work informs stronger evaluation standards that address central challenges in the current evaluation ecosystem. Our proposal has the added side benefit of creating incentives for model builders to create models that can be fine-tuned easily and respond well to fine-tuning.
@@ A.1 Models considered
| gemma-2b | 2024-02 | 2 | 2.0 | google/gemma-2b | Gemma et al. (2024) |
@@
| olmo-1.7-7b | 2024-04 | 7 | 2.05 | allenai/OLMo-1.7-7B-hf | Groeneveld et al. (2024) |
@@
| olmo-7b | 2024-01 | 7 | 2.46 | allenai/OLMo-7B-hf | Groeneveld et al. (2024) |
@@
| pythia-1.4b | 2023-02 | 1.4 | 0.3 | EleutherAI/pythia-1.4b | Biderman et al. (2023) |
| pythia-12b | 2023-02 | 12 | 0.3 | EleutherAI/pythia-12b | Biderman et al. (2023) |
| pythia-160m | 2023-02 | 0.16 | 0.3 | EleutherAI/pythia-160m | Biderman et al. (2023) |
| pythia-1b | 2023-02 | 1 | 0.3 | EleutherAI/pythia-1b | Biderman et al. (2023) |
| pythia-2.8b | 2023-02 | 2.8 | 0.3 | EleutherAI/pythia-2.8b | Biderman et al. (2023) |
| pythia-410m | 2023-02 | 0.41 | 0.3 | EleutherAI/pythia-410m | Biderman et al. (2023) |
| pythia-6.9b | 2023-02 | 6.9 | 0.3 | EleutherAI/pythia-6.9b | Biderman et al. (2023) |
| pythia-70m | 2023-02 | 0.07 | 0.3 | EleutherAI/pythia-70m | Biderman et al. (2023) |
@@ A.2 Fine-tuning hyperparameters
We fine-tune all model parameters. For models with less than B parameters, we fine-tune on a single GPU with BF16 precision. For models between B and B parameters, we train on a single H100 node using DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) and full precision. For models with more than B parameters, we train on two H100 nodes using DeepSpeed ZeRO-3 and full precision. Due to the large compute cost of the experiments, we perform minimal hyperparameter tuning and use standard hyperparameter choices throughout. We use a learning rate of for models with fewer than B parameters and a learning rate of for models with more than B parameters. For four of the B models –Gemma 7B, Olmo 7B, Olmo 1.7 7B, and Llama 3 8B– benchmark accuracy heavily degraded after fine-tuning. For these models, we use a peak learning rate of instead. These four models were all released after November 2023. We use a cosine learning rate schedule with linear warm-up for 50 steps and decay to of the peak learning rate. We use AdamW (Loshchilov and Hutter, 2018) as the optimizer, with , , and . We fine-tune with batch size 64. We use a weight decay rate of 0.1 and clip gradients at 1.0. We verify that the training loss decreases for all models on both of the fine-tuning datasets. To reduce the computation burden of fine-tuning, we train with context size 600. We verify that less than 5% of the fine-tuning examples have context length above 600.
@@ Appendix B Causal interpretation of our findings
In Section 2.2 we demonstrated that models trained after November 2023 significantly outperform those trained before November 2023 for both MMLU and GSM8K. We now seek to determine how much of the benchmark improvements of newer models is attributable to newer models training more on the test task. That is, the extent to which the effect of model recency on benchmark accuracy is mediated by training on the test task . The key obstacle to our analysis is that test task training is unobservable. Firstly, because practitioners are typically not transparent about their designs choices (e.g., pretraining data). Secondly, because the extent to which different training practices might amount to test task training is unclear. However, we are able to intervene on by fine-tuning on the test task.
Figure 10 summarizes our causal assumption. The time at which a model was trained determines the design choices made, such as its pretraining data or pretraining compute . These design choices in turn affect how much the model trains on the test task. All these factors ultimately influence the pretrained model and thus its benchmark performance. We assume that test task training does not causally influence pretraining compute, but compute might influence test task training. For instance, training on larger datasets may lead to training more on the test task.
We intervene on test task training by fine-tuning all models on the same, sufficient amount of task-specific data before evaluation. That is, we apply the adjustment proposed in Section 2.1. The external validity of our analysis hinges on the assumption that our experimental setting –fine-tuning models after the pretraining stage– is reasonably similar to the natural settings in which practitioners might train on the test task during pretraining (e.g., by including instruction data in the pretraining data mixture). We provide evidence in Appendix B.2 that this is the case.
We model fine-tuning as a hard intervention (Pearl, 2009). The specific magnitude of the intervention need not be quantified. Instead, the key assumption is that by fine-tuning on the same, sufficient amount of task data, all models will have received the same amount of test task training. Since some base models may have already trained on the test task prior to fine-tuning, our assumption will only hold if test task training saturates and we train on enough task data to reach saturation. We find evidence of saturation for both MMLU and GSM8K, see Appendix B.3.
@@
| MMLU | GSM8K | |
| 0.071 | 0.168 | |
| (0.018) | (0.032) | |
| R2 | 0.530 | 0.503 |
Standard errors in parentheses. Bold indicates .
We fit the models in Equation 5 and Equation 6, and we report the coefficients pertaining to and in Table 3 and Table 3, respectively. We find no evidence of a significant direct effect of model recency on accuracy. On the other hand, its indirect effect mediated by test task training is significant, positive, and large.
Therefore, our analysis indicates that the differences in MMLU and GSM8K performance between newer and older models observed in Section 2.1 are primarily attributable to differences in test task training. That is, the mechanism by which newer models outperform older models is by training more on the test task.
B.1 Robustness check on the temporal split: EN vs CN language data
@@
We repeat the analysis of Section 2 for the EN and EN+CN model split. We observe that, controlling for pretraining compute, models trained on EN+CN language data outperform those trained primarily on EN by 9 accuracy points on MMLU and 12 accuracy points on GSM8K. After the proposed adjustment, however, the difference in performance between models trained on EN data and EN+CN data is small and not statistically significant.
@@
B.2 How similar are newer models to older, fine-tuned models?
@@
We train XGBoost classifiers and report balanced accuracy for leave-one-out cross-validation in Table 4. We find that newer models are reasonable distinguishable from older models, with 63% accuracy for MMLU and 79% accuracy for GSM8K. In contrast, we obtain close to random-chance accuracy in discriminating between older, fine-tuned models and newer models. That is, older fine-tuned models are indistinguishable from newer models in terms of their performance.
B.3 Test task training and saturation
We show that training on the test task saturates. We consider the intermediate checkpoints of the adjustment procedure, that is, fine-tuning for 3 epochs of the task datasets introduced in Section 2.1. We plot in Figure 13 the gain in benchmark accuracy from the first 75% (MMLU) and 80% (GSM8K) training steps, as well as the gain in benchmark accuracy from the remaining steps. Almost all of the performance improvements occur in the earlier training steps. In contrast, the final 25% (MMLU) and 20% (GSM8K) optimization steps result in almost no changes in benchmark accuracy. This indicates that training on the test task saturates, and we train for enough steps to reach saturation.
Note that while more training data might result in further benchmark improvements, we show that the task datasets that we use are sufficient for older models to reach the performance of newer models, see Appendix B.2.
Appendix C Results for the OpenLLM Leaderboard v2
@@
We evaluate all 53 models on MMLU Pro, GPQA, BBH, MuSR and MATH Lvl 5. We use the LM Evaluation Harness library in identical fashion to the HF leaderboard v2. We do not evaluated on IFEval since it tests for instruction following and we evaluate base models. We additionally evaluate the models that we fine-tuned in Section 2.1 for multiple choice question answering and mathematical reasoning. This gives us models’ adjusted benchmark scores after training on multiple choice question answering and mathematical reasoning. For MATH Lvl 5, we use the models fine-tuned on mathematical data, whereas for MMLU Pro, GPQA, BBH and MuSR we use the models fine-tuned on multiple choice question answering. The fine-tuning datasets were not adapted to the new benchmarks in the HF leaderboard v2, thus giving a valuable insight into how well these task-relevant datasets generalize beyond MMLU and GSM8K.
@@
Fine-tuning on task-relevant data reduces the difference in performance between newer and older models for all five benchmarks. For GPQA and MuSR, the difference in performance after adjustment is small () and not statistically significant. For BBH, the estimated difference in performance reduces by 40% to 0.015 and is no longer statistically significant. For MMLU Pro and MATH Lvl 5 the difference reduces by 19% and 33% respectively but remains reasonably large ( > 0.01) and statistically significant. Therefore, find evidence that training on the test task plays a substantial role in newer models outperforming older ones in the benchmarks of the HF Leaderboard v2.
@@
One possible reason for the fact that the adjustment for MMLU Pro and MATH Lvl 5 is not as effective as for MMLU and GSM8K is that the fine-tuning examples are simply not as relevant for MMLU Pro and MATH Lvl 5. For example, the questions in MATH Lvl 5 contain much more LaTeX equation formatting than our mathematical reasoning fine-tuning dataset. Note that the answers to many MATH Lvl 5 questions are precisely formatted as LaTeX equations. Regarding MMLU Pro, our multiple choice fine-tuning dataset contains mostly questions with 4 answer choices, whereas all MMLU Pro questions have 10 answer choices. Thus, models are primarily fine-tuned to answer “A”, “B”, “C”, and “D” but not “E”, “F”, “G”. We modify MMLU Pro to only contain questions with 4 answer choices by for every question randomly discarding 6 of the incorrect answer choices. We evaluate models pre and post adjustment and plot the results in Figure 15. We observe that the difference in performance between newer and older models after adjustment reduces from 0.024 to 0.016, and is no longer statistically significant. This observation suggests that fine-tuning one more relevant task-data might further reduce the gap between newer and older models.

### v2

@@ 7 Discussion
The anecdote holds a lesson for the evaluation of large language models half a century later. Knowledge about the evaluation conditions necessarily influences training practices under competitive pressure. It may be a fool’s errand to prohibit the practice. Instead, we propose to adjust for it by giving every model the same task-specific preparation before evaluation. We work from the assumption that training on the test task, in general, cannot be effectively detected, disallowed, or disincentivized. Detecting what training data a model has seen is a notoriously difficult problem –existing heuristics achieve partial success at best. Researchers routinely acknowledge the futility of fighting data contamination. Moreover, we anticipate that the ways to effectively train on the test task will only grow in scope and adoption.
Our work demonstrates that comparisons of different models are confounded by the choice of training data and training practices. Different model families vary in the degree that they were—implicitly or explicitly—trained on various test tasks. A relatively small amount of task data can have a disproportional large effect on benchmark performance. It therefore makes little sense to compare model performance at face value without accounting for how the training data relate to the test task.
Training on the test task also has profound implications for the study of emergent behavior. After training on the test task, model capabilities become predictable at smaller scales. This greatly reduces the unpredictability associated with emergence, notably without any change in the metric.
Despite the daunting challenges that training on the test task poses for the fair evaluation of language models, it’s also its own best remedy. Giving each model the same sufficient task-specific fine-tuning harmonizes model comparisons and linearizes the relationship between model capabilities and pretraining compute. We hope that our work informs stronger evaluation standards that address central challenges in the current evaluation ecosystem.
@@ A.1 Models considered
| gemma-2b | 2024-02 | 2 | 3.0 | google/gemma-2b | Gemma et al. (2024) |
@@
| olmo-1.7-7b | 2024-04 | 7 | 2.0 | allenai/OLMo-1.7-7B-hf | Groeneveld et al. (2024) |
@@
| olmo-7b | 2024-01 | 7 | 2.5 | allenai/OLMo-7B-hf | Groeneveld et al. (2024) |
@@
| pythia-1.4b | 2022-10 | 1.4 | 0.3 | EleutherAI/pythia-1.4b | Biderman et al. (2023) |
| pythia-12b | 2022-10 | 12 | 0.3 | EleutherAI/pythia-12b | Biderman et al. (2023) |
| pythia-160m | 2022-10 | 0.16 | 0.3 | EleutherAI/pythia-160m | Biderman et al. (2023) |
| pythia-1b | 2022-10 | 1 | 0.3 | EleutherAI/pythia-1b | Biderman et al. (2023) |
| pythia-2.8b | 2022-10 | 2.8 | 0.3 | EleutherAI/pythia-2.8b | Biderman et al. (2023) |
| pythia-410m | 2022-10 | 0.41 | 0.3 | EleutherAI/pythia-410m | Biderman et al. (2023) |
| pythia-6.9b | 2022-10 | 6.9 | 0.3 | EleutherAI/pythia-6.9b | Biderman et al. (2023) |
| pythia-70m | 2022-10 | 0.07 | 0.3 | EleutherAI/pythia-70m | Biderman et al. (2023) |
@@
| qwen2-0.5b | 2024-06 | 0.5 | 12.0 | Qwen/Qwen2-0.5B | Yang et al. (2024) |
| qwen2-1.5b | 2024-06 | 1.5 | 7.0 | Qwen/Qwen2-1.5B | Yang et al. (2024) |
| qwen2-7b | 2024-06 | 7 | 7.0 | Qwen/Qwen2-7B | Yang et al. (2024) |
@@
A.1.1 Instruction-tuned and chat models
| Name | Train date | Base model | HF repository | Citation |
| falcon-7b-instruct | 2023-04 | falcon-7b | tiiuae/falcon-7b-instruct | Almazrouei et al. (2023) |
| gemma-2b-instruct | 2024-02 | gemma-2b | google/gemma-2b-it | Gemma et al. (2024) |
| gemma-7b-instruct | 2024-02 | gemma-7b | google/gemma-7b-it | Gemma et al. (2024) |
| internlm-chat-20b | 2023-09 | internlm-20b | internlm/internlm-chat-20b | InternLM (2023) |
| internlm-chat-7b | 2023-07 | internlm-7b | internlm/internlm-chat-7b | InternLM (2023) |
| internlm2-7b | 2024-01 | internlm2-base-7b | internlm/internlm2-7b | Cai et al. (2024) |
| internlm2-chat-1_8b | 2024-01 | internlm2-base-7b | internlm/internlm2-chat-1_8b | Cai et al. (2024) |
| internlm2-chat-20b | 2024-01 | internlm2-base-20b | internlm/internlm2-chat-20b | Cai et al. (2024) |
| internlm2-chat-7b | 2024-01 | internlm2-base-7b | internlm/internlm2-chat-7b | Cai et al. (2024) |
| llama-2-13b-chat | 2023-07 | llama-2-13b | meta-llama/Llama-2-13b-chat-hf | Touvron et al. (2023b) |
| llama-2-7b-chat | 2023-07 | llama-2-7b | meta-llama/Llama-2-7b-chat-hf | Touvron et al. (2023b) |
| llama-3-8b-instruct | 2024-04 | llama-3-8b | meta-llama/Meta-Llama-3-8B-Instruct | MetaAI (2024) |
| map-neo-7b-instruct | 2024-05 | map-neo-7b | m-a-p/neo_7b_instruct_v0.1 | Zhang et al. (2024) |
| map-neo-7b-sft | 2024-05 | map-neo-7b | m-a-p/neo_7b_sft_v0.1 | Zhang et al. (2024) |
| olmo-7b-0724-instruct-hf | 2024-01 | olmo-7b | allenai/OLMo-7B-0724-Instruct-hf | Groeneveld et al. (2024) |
| olmo-7b-0724-sft-hf | 2024-01 | olmo-7b | allenai/OLMo-7B-0724-SFT-hf | Groeneveld et al. (2024) |
| olmo-7b-instruct-hf | 2024-01 | olmo-7b | allenai/OLMo-7B-Instruct-hf | Groeneveld et al. (2024) |
| olmo-7b-sft-hf | 2024-01 | olmo-7b | allenai/OLMo-7B-SFT-hf | Groeneveld et al. (2024) |
| qwen-1.5-0.5b-chat | 2024-01 | qwen-1.5-0.5b | Qwen/Qwen1.5-0.5B-Chat | Bai et al. (2023) |
| qwen-1.5-1.8b-chat | 2024-01 | qwen-1.5-1.8b | Qwen/Qwen1.5-1.8B-Chat | Bai et al. (2023) |
| qwen-1.5-14b-chat | 2024-01 | qwen-1.5-14b | Qwen/Qwen1.5-14B-Chat | Bai et al. (2023) |
| qwen-1.5-4b-chat | 2024-01 | qwen-1.5-4b | Qwen/Qwen1.5-4B-Chat | Bai et al. (2023) |
| qwen-1.5-7b-chat | 2024-01 | qwen-1.5-7b | Qwen/Qwen1.5-7B-Chat | Bai et al. (2023) |
| redpajama-7b-chat | 2023-05 | redpajama-7b | togethercomputer/RedPajama-INCITE-7B-Chat | TogetherWeCompute (2023) |
| redpajama-chat-3b-v1 | 2023-05 | redpajama-3b | togethercomputer/RedPajama-INCITE-Chat-3B-v1 | TogetherWeCompute (2023) |
| redpajama-instruct-3b-v1 | 2023-05 | redpajama-3b | togethercomputer/RedPajama-INCITE-Instruct-3B-v1 | TogetherWeCompute (2023) |
| redpajama-instruct-7b | 2023-05 | redpajama-7b | togethercomputer/RedPajama-INCITE-7B-Instruct | TogetherWeCompute (2023) |
| stablelm-2-1.6b-chat | 2024-01 | stablelm-2-1.6b | stabilityai/stablelm-2-1_6b-chat | Bellagente et al. (2024) |
| stablelm-2-12b-chat | 2024-03 | stablelm-2-12b | stabilityai/stablelm-2-12b-chat | Bellagente et al. (2024) |
| stablelm-2-zephyr-1.6b | 2024-06 | stablelm-2-1.6b | stabilityai/stablelm-2-zephyr-1_6b | Bellagente et al. (2024) |
| stablelm-zephyr-3b | 2023-11 | stablelm-3b-4e1t | stabilityai/stablelm-zephyr-3b | Tunstall et al. (2023) |
| vicuna-13b-v1.1 | 2023-04 | llama-13b | lmsys/vicuna-13b-v1.1 | Chiang et al. (2023) |
| vicuna-13b-v1.3 | 2023-06 | llama-13b | lmsys/vicuna-13b-v1.3 | Chiang et al. (2023) |
| vicuna-7b-v1.1 | 2023-04 | llama-7b | lmsys/vicuna-7b-v1.1 | Chiang et al. (2023) |
| vicuna-7b-v1.3 | 2023-06 | llama-7b | lmsys/vicuna-7b-v1.3 | Chiang et al. (2023) |
@@ A.2 Fine-tuning hyperparameters
We fine-tune all model parameters. For models with less than B parameters, we fine-tune on a single GPU with BF16 precision. For models between B and B parameters, we train on a single H100 node using DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) and full precision. For models with more than B parameters, we train on two H100 nodes using DeepSpeed ZeRO-3 and full precision. Due to the large compute cost of the experiments, we perform minimal hyperparameter tuning and use standard hyperparameter choices throughout. We use a learning rate of for models with fewer than B parameters and a learning rate of for models with more than B parameters. For Qwen 2 as well as four of the B models –Gemma 7B, Olmo 7B, Olmo 1.7 7B, and Llama 3 8B– benchmark accuracy degraded after fine-tuning. For these models, we use a peak learning rate of instead. We use a cosine learning rate schedule with linear warm-up for 50 steps and decay to of the peak learning rate. We use AdamW (Loshchilov and Hutter, 2018) as the optimizer, with , , and . We fine-tune with batch size 64. We use a weight decay rate of 0.1 and clip gradients at 1.0. We verify that the training loss decreases for all models on both of the fine-tuning datasets. To reduce the computation burden of fine-tuning, we train with context size 600. We verify that less than 5% of the fine-tuning examples have context length above 600.
@@
A.3 Robustness check on the hyperparameter search budget
We found the learning rate to be, by a large margin, the single most impactful hyperparameter. We perform a sweep on the MMLU auxiliary training set with the following learning rates: [, , , , ]. We are unable to perform more extensive hyperparameter sweeps due to their large computational cost. We find that no models benefit from a smaller learning rate than , and only one model benefits from a larger learning rate than , Pythia 70M (the model with smallest pre-training compute). Thus, for every model except for Pythia 70M, the optimal learning rate is inside the boundary of the hyperparameter sweep.
A potential concern is that older models may appear to match the performance of newer models because the fine-tuning hyperparameters could be systematically more favorable to the older models. To address this concern, we recreate our main experiment by fine-tuning older models with their original learning rates while selecting for the newer models the optimal learning rate in the sweep that leads to highest MMLU performance. That is, we deliberately give newer models a systematic advantage in terms of hyperparameter search budget. We plot the results in Figure 9. The estimated effect size of model recency on benchmark performance shifts from to . Therefore, despite giving newer models a systemic advantage in terms of hyperparameter search budget, the effect size remains both remains small and not statistically significant.
@@ Appendix B Causal interpretation of our findings
In Section 2.2 we established that models trained after November 2023 significantly outperform those trained before November 2023 for both MMLU and GSM8K. We then showed that fine-tuning all models in the test task equalizes the performance of newer and older models. We now present a causal interpretation of our findings, aiming to quantify the extent to which the effect of model recency on benchmark accuracy is mediated by training on the test task .
The key obstacle to our analysis is that test task training is unobservable. Firstly, because practitioners are typically not transparent about their designs choices, including the pretraining data. Secondly, because the extent to which different training practices might amount to test task training is unclear. Nonetheless, by fine-tuning on task-specific data we are able to intervene on the extent to which models train on the test task.
Figure 10 summarizes our causal assumption. The time at which a model was trained determines the design choices made, such as its pretraining data or pretraining compute . These design choices in turn affect how much the model trains on the test task. All these factors ultimately influence the pretrained model and thus its benchmark performance. We also admit that compute might influence test task training. For instance, pre-training on larger datasets may lead to models training more on the test task.
We interpret the proposed adjustment method as intervening on the test task training variable . Namely, by fine-tuning all models on the same amount of task-specific data before evaluation. The external validity of our subsequent analysis hinges on the assumption that our controlled experimental setting –fine-tuning models after the pretraining stage– is reasonably similar to the natural settings in which practitioners might train on the test task during pretraining (e.g., by including instruction data in the pretraining data mixture). We provide evidence for this in Appendix C.3.
We model fine-tuning as a hard intervention (Pearl, 2009). The specific magnitude of the intervention need not be quantified. Instead, the key assumption is that by fine-tuning on the same, sufficient amount of task data, all models will have received the same amount of test task training. Since some base models may have already trained on the test task prior to fine-tuning, this assumption only holds if test task training saturates and we train on enough task data to reach saturation. The fact that our task-specific datasets allow older models to match the performance of newer models provides some evidence that we train on enough task-specific data to reach saturation.
@@
| 0.071 | 0.168 | |
| (0.018) | (0.032) | |
| R2 | 0.530 | 0.503 |
Standard errors in parentheses. Bold indicates .
| MMLU | GSM8K | |
@@
We fit the models in Equation 5 and Equation 6, and we report the coefficients pertaining to and in Table 4 and Table 4. We find that the indirect effect of model recency on accuracy mediated by test task training is significant, positive, and large. In contrast, we find no evidence of a significant direct effect of model recency on accuracy. We therefore find no evidence of the improvements of newer models being attributable to anything else other than training on the test task.
In conclusion, our causal analysis indicates that the differences in MMLU and GSM8K performance between newer and older models observed in Section 2.1 are largely attributable to differences in test task training. That is, the mechanism by which newer models outperform older models is primarily by training more on the test task.
Appendix C Robustness check on the temporal split
C.1 Adjusting the temporal cutoff by a few months
We repeat the analysis of Section 2 for two additional temporal splits: September 2023 and January 2024, and present the results in Figure 11 and Figure 12, respectively. Our results are robust to shifting the temporal cutoff by a few months. That is, our findings indicate that practitioners started adopting design choices around late 2023 that amount to models training on the test task much more, which is consistent with models’ technical reports starting to mention the use of benchmark or instruction data at pre-training time. Choosing specifically the month of November as the cut-off is therefore not critical for our analysis.
C.2 EN vs CN language data
@@
We repeat the analysis of Section 2 for the EN and EN+CN model split, see Figure 13. We observe that, controlling for pretraining compute, models trained on EN+CN language data outperform those trained primarily on EN by 9 accuracy points on MMLU and 12 accuracy points on GSM8K. After the proposed adjustment, however, the difference in performance between models trained on EN data and EN+CN data is small and not statistically significant.
@@
C.3 How similar are newer models to older, fine-tuned models?
@@
We train XGBoost classifiers and report balanced accuracy for leave-one-out cross-validation in Table 5. We obtain close to random-chance accuracy in discriminating between older, fine-tuned models and newer models. That is, older fine-tuned models are indistinguishable from newer models in terms of their performance per pre-training compute.
Appendix D Results for the OpenLLM Leaderboard v2
@@
We evaluate all models on MMLU Pro, GPQA, BBH, MuSR and MATH Lvl 5. We use the LM Evaluation Harness library in identical fashion to the HF leaderboard v2. We do not evaluated on IFEval since it tests for instruction following and we evaluate base models. We additionally evaluate the models that we fine-tuned in Section 2.1 for multiple choice question answering and mathematical reasoning. This gives us models’ adjusted benchmark scores after training on multiple choice question answering and mathematical reasoning. For MATH Lvl 5, we use the models fine-tuned on mathematical data, whereas for MMLU Pro, GPQA, BBH and MuSR we use the models fine-tuned on multiple choice question answering. The fine-tuning datasets were not adapted to the new benchmarks in the HF leaderboard v2, thus giving a valuable insight into how well these task-relevant datasets generalize beyond MMLU and GSM8K.
@@
Fine-tuning on task-relevant data reduces the difference in performance between newer and older models for all five benchmarks. Therefore, we find evidence that training on the test task plays a substantial role in newer models outperforming older ones in the benchmarks of the HF Leaderboard v2. For GPQA and MuSR, the difference in performance after adjustment is very small () and not statistically significant. For BBH, the estimated difference in performance reduces by 40% to 0.015 and is no longer statistically significant. For MMLU Pro and MATH Lvl 5 the difference reduces by 19% and 33% respectively but remains reasonably large ( > 0.01).
@@
One possible reason for the fact that the adjustment for MMLU Pro and MATH Lvl 5 is not as effective as for MMLU and GSM8K is that the fine-tuning examples are simply not as relevant for MMLU Pro and MATH Lvl 5. For example, the questions and answers in MATH Lvl 5 contain much more LaTeX equation formatting than our mathematical reasoning fine-tuning dataset. Similarly, our multiple choice fine-tuning dataset contains mostly questions with 4 answer choices, whereas all MMLU Pro questions have 10 answer choices. Thus, models are primarily fine-tuned to answer “A”, “B”, “C”, and “D” but not “E”, “F”, “G”. We modify MMLU Pro to contain questions with 4 answer choices by randomly discarding 6 of the incorrect answer choices. We evaluate models pre and post adjustment and plot the results in Figure 16. We observe that the difference in performance between newer and older models after adjustment reduces from 0.024 to 0.016, and is no longer statistically significant. This observation suggests that fine-tuning one more relevant task-data might further reduce the gap between newer and older models in MMLU Pro and MATH Lvl 5.